# Hypergraph Characterization of Fusion Rings

### v1

@@ 4.2. Simple Components
Suppose not. Then there exists a vertex $v$ which has a neighbor $u$ with degree at most 1 in $C[N(v)]$. In fact, $u$ must have a exactly one neighbor in $C[N(v)]$ since otherwise, by Equation 1, $\{u,v\}$ would participate in $-1$ hyperedges. Let $x$ be the unique vertex that is adjacent to both $u$ and $v$. Then noting that $w_{uv}=0$ by Equation 1, we have that
@@
If $\mathcal{C}$ is a multiplicity-free, self-dual, braided fusion category such that $N_{i,j}^{j}=N_{j,i}^{i}$ for all $i,j\neq 0$ and $N_{i,i}^{i}=0$, then for all non-invertible $X\in\operatorname{Irr}\left(\mathcal{C}\right)$, $X\otimes X$ contains at least three non-isomorphic subobjects.
@@
Let $H$ be a hypergraph such that $(G,H)$ generates a fusion ring and let $C$ be a non-trivial simple component of $G$ with minimum degree at most $3$. As a consequence of Lemma 1, the component $C$ contains a vertex $\ell$ of degree exactly 3. Furthermore, $C[N(\ell)\cup\{\ell\}]$ is a complete graph on 4 vertices, $\{i,j,k,\ell\}.$ By Equation 1 we have that $w_{\ell i}=w_{\ell j}=w_{\ell k}=1$. Suppose now that $\set{i,j,\ell}$ is the unique hyperedge involving $i$ and $\ell$, and note that
@@
which is a contradiction to the commutativity of $N_{\ell}$ and $N_{i}$. As $i$ and $j$ were arbitrary, we have that $\set{i,k,\ell}$ and $\set{j,k,\ell}$ are also not hyperedges in $H$.
Thus, there is some vertex $s\notin\set{i,j,k,\ell}$ such that $\set{s,i,\ell}$ is a hyperedge. Now note that
@@
As s≁not-$s\not\sim\ell$ and $w_{\ell s}\geq 3$, we have by Equation 1 that $\ell$ and $s$ have at least two common neighbors in $G$. Without loss of generality suppose that $i$ and $j$ are common neighbors of $\ell$ and $s$. At this point, we have that either $s\not\sim k$ and $w_{\ell s}=3$ or $s\sim k$ and there exists a vertex $t\not\in\{i,j,k\}$ such that $\{\ell,s,t\}\in E(H)$. Now considering
@@
we have that $1+e_{ks}=2+N_{\ell s}^{t}e_{k,t}$. Thus, $s\sim k$, $\{\ell,s,t\}\in E(H)$, and $k\sim t$. But as we now know that $k\sim s$, this implies that $i$ and $j$ were chosen arbitrarily. Repeating the previous argument for $\set{i,k}$ and $\set{j,k}$ and using that $w_{\ell s}=4$, we get that $i\sim t$ and $j\sim t$. As a result we have the edges and hyperedges depicted in Figure 3.
$\ell$$i$$j$$k$$s$$t$
$\ell$$i$$j$$k$$s$$t$
@@
Corollary 1.
If $\mathcal{C}$ is a self-dual, multiplicity-free, braided fusion category then either there exists a non-trivial simple object $X$ such that $\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)\ncong 0$, or for all non-invertible $Y\in\operatorname{Irr}\mathcal{C}$ $Y\otimes Y$ contains at least 4 non-isomorphic non-trivial subobjects.
Remark 1.
@@ 4.3. Non-simple Components
If $\mathcal{C}$ is a multiplcity-free, self-dual, braided fusion category, then for $X\ncong Y\in\operatorname{Irr}\left(\mathcal{C}\right)$ either $Y$ is a subobject of $X\otimes X$, or $X\otimes X$ and $Y\otimes Y$ have a common simple subobject.
@@
Let $G$ be an undirected graph that generates a fusion ring. If $G$ contains an induced path, $i\sim j\sim k$ where $\set{i,j}$ is not in a triangle, then $i$ and $j$ both have self loops.
@@
Let $\mathcal{C}$ be a multiplicity-free, self-dual, braided fusion category. If $X,Y,Z$ are pairwise non-isomorphic simple objects such that $Y$ is a subobject of $X\otimes X$ and $Z$ is a subobject of $Y\otimes Y$, but $X\otimes X$ and $Y\otimes Y$ have no common subobjects other than potentially $X$ and $Y$, then $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,Y\right)=1$.
@@ 4.4. Fusion Rings Generated by Triangle-Free Graphs
    a collection of $12^{k}-$ 1 isolated loopless vertices.
@@
Let $G$ be a triangle-free, undirected graph which generates a fusion ring. In the case that $G$ is the empty graph, by Theorem 1, $G$ has order $12^{k}-$ 1 and has a unique hypergraph $H$ associated with which it generates a fusion ring. Thus we may assume that $G$ contains at least one edge and let $T$ be the maximal collection of nontrivial triangle-free components. By Lemma 3 every edge in $T$ has a self-loop on at least one of the vertices. But then by Lemma 4 the diameter of $T$ is at most $4$ and in particular $T$ is a single component. In particular, $G$ consists of $T$ and some number of trivial (that is, isolated without a self-loop) vertices.
Suppose now that $T$ contains at least 3 vertices. Since $T$ is connected and contains at least $3$ vertices, for any edge $\set{i,j}$ in $T$, there exists some vertex $k$ such that either $i\sim k$ or $j\sim k$. Since $T$ is triangle-free, by Lemma 5 both $i$ and $j$ have a self-loop. As $\set{i,j}$ is an arbitrary edge, this implies that all vertices in $T$ have a self-loop and by Lemma 4 $T$ has diameter at most 2. In fact, since $T$ is triangle-free and has at least three vertices, this implies that the diameter of $T$ is exactly 2. Furthermore, by Lemma 3 for every edge $\set{i,j}\in E(T)$, $w_{ij}=1$. In particular, any hypergraph which, with $G$, generates a fusion ring is non-empty. Let $H$ be one such hypergraph and let $f=\{i,j,k\}$ be an arbitrary edge in $H$. Since $T$ is triangle-free, there are at most two edges of $T$ in $G[f]$. Suppose then that there are precisely two edges from $T$ in $f$, namely $i\sim j$ and $j\sim k$. Since $i,j,k$ are all in $T$, we have that $e_{ii}=e_{jj}=e_{kk}=1$ and, by Lemma 3, $w_{ik}=w_{kj}=1$. Furthermore, the hyperedge indicated by $w_{ij}$ and $w_{jk}$ is $f$. But now we have
@@
First suppose that $T$ is a looped vertex $x$. By Lemma 3 any trivial vertices must occur in exactly one hyperedge with $x$. This would require two isolated vertices, $i$ and $j$, being contained in a hyperedge with a looped vertex. However, this implies that $01=(N_{i}N_{j})_{ix}=(N_{i}N_{j})_{xi}=$ 0, a contradiction. Thus, if $T$ is a single looped vertex, then $T$ is the entire graph $G$. Next, suppose that $T$ is an edge $\{x,y\}$ with at least one loop. Without loss of generality, first suppose that only $x$ is looped. Since $w_{xy}=0$, additional trivial vertices must occur, in pairs, in exactly one hyperedge with $x$. Again we have that $T$ is the entire graph $G$. Finally suppose that both $x$ and $y$ are looped. By Lemma 3, $w_{xy}=1$ and thus at least one trivial vertex must be present. However, if more than one trivial vertex is present, a hyperedge of the form $\{x,i,j\}$ must also be present, where $i$ and $j$ are trivial vertices. Thus, such an edge must occur with exactly one trivial vertices. ∎
@@ 5. Low Rank Fusion Rings
If two digraphs, $D_{1}$ and $D_{2}$, generate fusion rings $F_{1}$ and $F_{2}$, there is a digraph $D$ that is a super graph of $D_{1}\times D_{2}$ which generates $F_{1}\boxtimes F_{2}$. More concretely let $D_{1}^{\prime}$ be formed from $D_{1}$ by adding the vertex 00 which has a self-loop and directed edges from every vertex in $D_{1}$ to 00. Then $D=D_{1}^{\prime}\times D_{2}^{\prime}-\set{(0,0)}$ and $D$ generates the fusion ring $F_{1}\boxtimes F_{2}$.
Remark 2.
@@
$\left(\sum_{i}N_{ii}^{i},\sum_{i}\sum_{j\neq i}N_{ii}^{j},\sum_{\left|\set{i,j,k}\right|=3}N_{ij}^{k}\right)$
uniquely determines the fusion ring. Furthermore, in examining the rank 5 fusion rings we see that the more algebraically natural tuples $\left(\sum_{i}\operatorname{Tr}\!\left(N_{i}\right),\sum_{\left|\set{i,j,k}\right|=3}N_{ij}^{k}\right)$ or $\left(\sum_{i}\operatorname{Tr}\!\left(N_{i}\right),\sum_{i,j,k}N_{i,j}^{k}\right)$ fail to distinguish between all fusion rings. We are not so bold as to conjecture that this holds in general, but it may be a fruitful direction of future research.
@@
$1$ $(1,2)$ $\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[20]
$1$ $(1,2)$, $(2,1)$ $\operatorname{PSU}(3)_{2}$
@@
$2$ $(1,2)$ $(1,2,3)$ $\operatorname{Sem}\boxtimes\operatorname{Fib}$[21]
$1$, $2$ $(1,2)$, $(2,1)$ $(1,2,3)$ $\operatorname{PSU}(2)_{6}$[6]
$1$, $2$ $(1,2)$, $(1,3)$, $(2,1)$, $(3,2)$ $(1,2,3)$ $\operatorname{PSU}(2)_{5}$[6]
$1$, $2$, $3$ $(1,2)$, $(1,3)$ $(1,2,3)$ $\operatorname{Fib}^{2}$ [21]
@@
$1$ $(1,2)$, $(1,3)$, $(1,4)$ $(2,3,4)$ not categorifiable[9]
$1$ $(1,4)$, $(2,1)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$ $\mathcal{C}(\operatorname{\mathfrak{so}}_{3},e^{\nicefrac{{\pi i}}{{6}}},6)$
$1$, $3$ $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$
$1$, $3$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$
$2$, $3$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(3,2)$ $(1,2,3)$, $(2,3,4)$
$1$, $2$, $3$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(2,3,4)$
$1$, $2$, $3$ $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(3,2)$ $(1,2,3)$, $(2,3,4)$
$1$, $2$, $3$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,4)$, $(3,2)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$

### v2

@@ 4.2. Simple Components
Suppose not. Then there exists a vertex $v$ which has a neighbor $u$ with degree at most 1 in $C[N(v)]$. In fact, $u$ must have a exactly one neighbor in $C[N(v)]$ since otherwise, by Equation 1, $\{u,v\}$ would participate in 1-1 hyperedges. Let $x$ be the unique vertex that is adjacent to both $u$ and $v$. Then noting that $w_{uv}=0$ by Equation 1, we have that
@@
If $\mathcal{C}$ is a multiplicity-free, self-dual, braided fusion category such that $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)=0$ for all $X\in\operatorname{Irr}^{\prime}\!(\mathcal{C})$ and $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,Y\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,X\right)$ for all distinct objects $X,Y\in\operatorname{Irr}^{\prime}\!(\mathcal{C})$, then then for all non-invertible $\in\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right), XX$$\otimes X$ contains at least three non-isomorphic subobjects from $\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$.
@@
Let $H$ be a hypergraph such that $(G,H)$ generates a fusion ring and let $C$ be a non-trivial simple component of $G$ with minimum degree at most $3$. As a consequence of Lemma 1, the component $C$ contains a vertex $\ell$ of degree exactly 3. Furthermore, $C[N(\ell)\cup\{\ell\}]$ is a complete graph on 4 vertices, $\{i,j,k,\ell\}.$ By Equation 1 we have that $w_{\ell i}=w_{\ell j}=w_{\ell k}=1$. Suppose now that $\{i,j,\ell\}$ is the unique hyperedge involving $i$ and $\ell$, and note that
@@
which is a contradiction to the commutativity of $N_{\ell}$ and $N_{i}$. As $i$ and $j$ were arbitrary, we have that $\{i,k,\ell\}$ and $\{j,k,\ell\}$ are also not hyperedges in $H$.
Thus, there is some vertex $s\notin\{i,j,k,\ell\}$ such that $\{s,i,\ell\}$ is a hyperedge. Now note that
@@
As $s\not\sim\ell$ and $w_{\ell s}\geq 3$, we have by Equation 1 that $\ell$ and $s$ have at least two common neighbors in $G$. Without loss of generality suppose that $i$ and $j$ are common neighbors of $\ell$ and $s$. At this point, we have that either $s\not\sim k$ and $w_{\ell s}=3$ or $s\sim k$ and there exists a vertex $t\not\in\{i,j,k\}$ such that $\{\ell,s,t\}\in E(H)$. Now considering
@@
we have that $1+e_{ks}=2+N_{\ell s}^{t}e_{k,t}$. Thus, $s\sim k$, $\{\ell,s,t\}\in E(H)$, and $k\sim t$. But as we now know that $k\sim s$, this implies that $i$ and $j$ were chosen arbitrarily. Repeating the previous argument for $\{i,k\}$ and $\{j,k\}$ and using that $w_{\ell s}=4$, we get that $i\sim t$ and $j\sim t$. As a result we have the edges and hyperedges depicted in Figure 3.
[…]
@@
Algebraic Corollary 3.
If $\mathcal{C}$ is a self-dual, multiplicity-free, braided fusion category such that $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)=0$ for all $X\in\operatorname{Irr}^{\prime}\!(\mathcal{C})$ and $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,Y\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,X\right)$ for all distinct objects $X,Y\in\operatorname{Irr}^{\prime}\!(\mathcal{C})$, then all for all non-invertible $Y\in\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$ the product $Y\otimes Y$ contains at least 4 non-isomorphic subobjects from $\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$.
Remark.
@@ 4.3. Non-simple Components
If $\mathcal{C}$ is a multiplcity-free, self-dual, braided fusion category such that $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,Y\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,X\right)$ for all distinct objects $X,Y\in\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$, then for distinct subobjects $X,Y\in\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$ such that $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,Y\right)=1$ either $Y$ is a subobject of $X\otimes X$, or $X\otimes X$ and $Y\otimes Y$ have a common subobject in $\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$.
@@
Let $G$ be an undirected graph that generates a fusion ring. If $G$ contains an induced path, $i\sim j\sim k$ where $\{i,j\}$ is not in a triangle, then $i$ and $j$ both have self loops.
@@
Let $\mathcal{C}$ be a multiplicity-free, self-dual, braided fusion category such that $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,Y\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,X\right)$ for all distinct objects $X,Y\in\operatorname{Irr}^{\prime}\!(\mathcal{C})$. If $X,Y,Z$ are distinct elements of $\operatorname{Irr}^{\prime}\!\left(\mathcal{C}\right)$ such that $X$ and $Z$ are subobjects of $Y\otimes Y$ and $Z$ is not a subobject of $X\otimes X$, then either, $X\otimes X$ and $Y\otimes Y$ have a common subobject other than $\mathbbm{1}$, $X$, and $Y$ or $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,Y\right)=1$.
@@ 4.4. Fusion Rings Generated by Triangle-Free Graphs
    a collection of $2^{k}-1$ isolated loopless vertices.
@@
Let $G$ be a triangle-free, undirected graph which generates a fusion ring. In the case that $G$ is the empty graph, by Theorem 1, $G$ has order $2^{k}-1$ and has a unique hypergraph $H$ associated with which it generates a fusion ring. Thus we may assume that $G$ contains at least one edge and let $T$ be the maximal collection of nontrivial triangle-free components. By Lemma 3 every edge in $T$ has a self-loop on at least one of the vertices. But then by Lemma 4 the diameter of $T$ is at most $4$ and in particular $T$ is a single component. In particular, $G$ consists of $T$ and some number of trivial (that is, isolated without a self-loop) vertices.
Suppose now that $T$ contains at least 3 vertices. Since $T$ is connected and contains at least $3$ vertices, for any edge $\{i,j\}$ in $T$, there exists some vertex $k$ such that either $i\sim k$ or $j\sim k$. Since $T$ is triangle-free, by Lemma 5 both $i$ and $j$ have a self-loop. As $\{i,j\}$ is an arbitrary edge, this implies that all vertices in $T$ have a self-loop and by Lemma 4 $T$ has diameter at most 2. In fact, since $T$ is triangle-free and has at least three vertices, this implies that the diameter of $T$ is exactly 2. Furthermore, by Lemma 3 for every edge $\{i,j\}\in E(T)$, $w_{ij}=1$. In particular, any hypergraph which, with $G$, generates a fusion ring is non-empty. Let $H$ be one such hypergraph and let $f=\{i,j,k\}$ be an arbitrary edge in $H$. Since $T$ is triangle-free, there are at most two edges of $T$ in $G[f]$. Suppose then that there are precisely two edges from $T$ in $f$, namely $i\sim j$ and $j\sim k$. Since $i,j,k$ are all in $T$, we have that $e_{ii}=e_{jj}=e_{kk}=1$ and, by Lemma 3, $w_{ik}=w_{kj}=1$. Furthermore, the hyperedge indicated by $w_{ij}$ and $w_{jk}$ is $f$. But now we have
@@
First suppose that $T$ is a looped vertex $x$. By Lemma 3 any trivial vertices must occur in exactly one hyperedge with $x$. This would require two isolated vertices, $i$ and $j$, being contained in a hyperedge with a looped vertex. However, this implies that $1=(N_{i}N_{j})_{ix}=(N_{i}N_{j})_{xi}=0$, a contradiction. Thus, if $T$ is a single looped vertex, then $T$ is the entire graph $G$. Next, suppose that $T$ is an edge $\{x,y\}$ with at least one loop. Without loss of generality, first suppose that only $x$ is looped. Since $w_{xy}=0$, additional trivial vertices must occur, in pairs, in exactly one hyperedge with $x$. Again we have that $T$ is the entire graph $G$. Finally suppose that both $x$ and $y$ are looped. By Lemma 3, $w_{xy}=1$ and thus at least one trivial vertex must be present. However, if more than one trivial vertex is present, a hyperedge of the form $\{x,i,j\}$ must also be present, where $i$ and $j$ are trivial vertices. Thus, such an edge must occur with exactly one trivial vertices. ∎
@@ 5. Low Rank Fusion Rings
If two digraphs, $D_{1}$ and $D_{2}$, generate fusion rings $F_{1}$ and $F_{2}$, there is a digraph $D$ that is a super graph of $D_{1}\times D_{2}$ which generates $F_{1}\boxtimes F_{2}$. More concretely let $D_{1}^{\prime}$ be formed from $D_{1}$ by adding the vertex 0 which has a self-loop and directed edges from every vertex in $D_{1}$ to 0. Then $D=D_{1}^{\prime}\times D_{2}^{\prime}-\{(0,0)\}$ and $D$ generates the fusion ring $F_{1}\boxtimes F_{2}$.
Remark.
@@
$\left(\sum_{i}N_{ii}^{i},\sum_{i}\sum_{j\neq i}N_{ii}^{j},\sum_{\left|\{i,j,k\}\right|=3}N_{ij}^{k}\right)$
uniquely determines the fusion ring. Furthermore, in examining the rank 5 fusion rings we see that the more algebraically natural tuples $\left(\sum_{i}\operatorname{Tr}\!\left(N_{i}\right),\sum_{\left|\{i,j,k\}\right|=3}N_{ij}^{k}\right)$ or $\left(\sum_{i}\operatorname{Tr}\!\left(N_{i}\right),\sum_{i,j,k}N_{i,j}^{k}\right)$ fail to distinguish between all fusion rings. We are not so bold as to conjecture that this holds in general, but it may be a fruitful direction of future research.
Acknowledgments
The last author would like to thank Tobias Hagge for his help in formulating the algebraic corollaries as well as Gert Vercleyen and Sébastien Palcoux for their helpful pointers to more recent results on the enumeration of fusion rings and categories.
@@ 11 $\operatorname{Fib}$[24]
11 $(1,2)$ $\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[23]
11 $(1,2)$, $(2,1)$ $\operatorname{PSU}(3)_{2}$
@@
22 $(1,2)$ $(1,2,3)$ $\operatorname{Sem}\boxtimes\operatorname{Fib}$[24]
11, 22 $(1,2)$, $(2,1)$ $(1,2,3)$ $\operatorname{PSU}(2)_{6}$[8]
11, 22 $(1,2)$, $(1,3)$, $(2,1)$, $(3,2)$ $(1,2,3)$ $\operatorname{PSU}(2)_{5}$[8]
11, 22, 33 $(1,2)$, $(1,3)$ $(1,2,3)$ $\operatorname{Fib}^{2}$ [24]
@@ 22 $(1,2)$ $(1,2,3)$ $\operatorname{Sem}\boxtimes\operatorname{Fib}$[24]
11 $(1,2)$, $(1,3)$, $(1,4)$ $(2,3,4)$ not categorifiable[11]
11 $(1,4)$, $(2,1)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$ $\mathcal{C}(\operatorname{\mathfrak{so}}_{3},e^{\nicefrac{{\pi i}}{{6}}},6)$
11, 33 $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$
11, 33 $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$
22, 33 $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(3,2)$ $(1,2,3)$, $(2,3,4)$
11, 22, 33 $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(2,3,4)$
11, 22, 33 $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(3,2)$ $(1,2,3)$, $(2,3,4)$
11, 22, 33, 44 $(1,2)$, $(1,3)$, $(1,4)$, $(2,4)$, $(3,2)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$